### v1

# Attosecond physics hidden in Cherenkov radiation

Abstract
Cherenkov radiation of charged particles moving with superluminal velocities in transparent media is a well-studied phenomenon with a plethora of applications. Its microscopic origins can be traced to the polarization of atomic shells, characterized by time scales in the subfemtosecond range — dynamics that eludes conventional macroscopic treatment. Here we present a theoretical framework for probing the intrinsic dynamics of Cherenkov radiation, unveiling quantum features absent in classical realm and even in a fully quantum theory in momentum space. These features include a finite formation length and spreading time of the photon, the latter becoming negative nearby the Cherenkov angle, a finite flash duration tied to the size of the electron packet, along with a shift in the photon arrival time that can be either positive or negative. The calculated time scales naturally lie in the attosecond range for relevant parameters, thereby linking this macroscopic phenomenon back to its atomic origins. Finally, we propose that by measuring the duration of the Cherenkov flash one can in principle retrieve the length of the emitting packet, deepening our understanding of quantum coherence effects in photon emission.
Introduction
Cherenkov radiation (ChR) by charged particles in media [1, 2, 3, 4, 5, 6, 7, 8, 9] is the simplest example of a wide range of phenomena embracing transition, diffraction, Smith-Purcell, and other mechanisms of photon emission [10, 11, 12, 13]. Their common microscopic origin is atomic bremsstrahlung [14] due to dynamic polarization of atomic shells by the field of the charge, and the characteristic time scales are femto- and attoseconds, typical for the AC Stark effect [15]. Despite the growing interest to quantum features of ChR [16, 17, 18], a solid link between microscopic and macroscopic theories remains elusive, and the atomic dynamics stays hidden even in the fully quantum treatment of ChR in momentum space and of its generalizations.
Here we point out that one can access atomic time scales in ChR by using a quantum theory in phase space. Similar to quantum optics [19], we employ a Wigner function to characterize the emitted photon, which depends on a momentum ${\bf k}$, on a point ${\bf r}$ in space and a time instant $t$. We demonstrate how to probe the field in the pre-wave (or formation) zone [20, 11] in which the partial waves interfere, spread, and the Cherenkov cone is not formed yet. Most strikingly, our theory establishes a link between the evolving coherence length of the electron packet, the Cherenkov flash duration, and a quantum temporal delay that the photon experiences in medium. This delay can be either positive or negative and belongs to the attosecond range, typical for atomic excitation processes studied by the attosecond spectroscopy and metrology [21]. The quantum shift in the photon arrival time, coupled with the finite flash duration and the spreading dynamics, unveils intricate coherence properties that enable refined temporal control in quantum emission processes, enriching the landscape of quantum optics, ultrafast physics, and of applications of ChR and related phenomena. The system of units $\hbar=c=1$ is used, the electron velocity is $\beta=u_{p}/c\equiv u_{p}$, $m=0.511$ MeV is the electron mass, $1/m\equiv\hbar/mc=3.86\times 10^{-11}$ cm, and $t_{c}=1/m\equiv\hbar/mc^{2}\approx 1.3\times 10^{-21}\,\text{s}$.
Figure 1: Generation of Cherenkov radiation by an electron packet with an rms size $\sigma_{x}$ in a dielectric medium with a refractive index $n$. The photon field gradually becomes separated from the electron in the formation zone and spreads. The Cherenkov cone is formed in the far-field only if the electron is not detected. The phase space picture implies that the photon with a momentum ${\bf k}$ is detected in a region of space centered at the point ${\bf r}$ at the moment of time $t$.
Photon emission in phase space
When an electron emits a photon, the two-particle state within the first order of the perturbation theory is $|e^{\prime},\gamma\rangle=(\hat{1}+\hat{S}^{(1)})|\text{in}\rangle$, where $|\text{in}\rangle=|e_{\text{in}}\rangle\otimes|0_{\gamma}\rangle$ and $\hat{S}^{(1)}=-ie\int d^{4}x\,\hat{j}^{\mu}(x)\hat{A}_{\mu}(x)$ [22]. Fig.1 illustrates how the photon packet gradually becomes spatially separated from the electron, spreads, and eventually turns into a plane wave propagating at the angles $\theta,\phi$. If we detect the electron in a plane-wave state $\langle{\bf p}^{\prime},\lambda^{\prime}|$, $\lambda^{\prime}=\pm 1/2$, the state of the photon becomes
$\displaystyle|\gamma\rangle=\langle{\bf p}^{\prime},\lambda^{\prime}|e_{\text{in}}\rangle|0_{\gamma}\rangle+\sum\limits_{\lambda_{\gamma}}\int\frac{d^{3}k}{(2\pi)^{3}}\,|{\bf k},\lambda_{\gamma}\rangle\,S_{fi}.$ (1)
The second term here is a coherent superposition of plane waves with the momenta ${\bf k}$ and the helicity $\lambda_{\gamma}=\pm 1$ and a matrix element is $S_{fi}=\langle{\bf k},\lambda_{\gamma};{\bf p}^{\prime},\lambda^{\prime}|\hat{S}^{(1)}|\text{in}\rangle$. We treat the incoming electron as a Gaussian packet with a wave function $f_{e}^{(\text{in})}({\bf p},\lambda)=\langle{\bf p},\lambda|e_{\text{in}}\rangle$, a mean momentum $\langle{\bf p}\rangle$, an uncertainty $\sigma\ll m$, which means that the rms size – also called coherence length – of the packet is much larger than the Compton wavelength, $\sigma_{x}=1/\sigma\gg 1/m$. For the moment, we take a simplified model with the packet, spherically symmetric in the laboratory frame, $\sigma_{\perp}=\sigma_{||}\equiv\sigma_{x}=1/\sigma$.
Let us define the energy density of the photon field in real space and time, an electric part of which is (Sec.1,2 in Supplementary)
$\displaystyle\frac{{\displaystyle 1}}{{\displaystyle 8\pi}}\langle\gamma|\hat{\bm{E}}^{2}({\bm{r}},t)|\gamma\rangle\to\frac{{\displaystyle 1}}{{\displaystyle 4\pi}}\left|\langle 0|\hat{\bm{E}}({\bm{r}},t)|\gamma\rangle\right|^{2}=\int\frac{d^{3}k}{(2\pi)^{3}}\,\mathcal{W}({\bf r},{\bf k},t),$ (2)
where we have subtracted a contribution of the vacuum energy and $\hat{\bm{E}}({\bm{r}},t)$ is a secondary-quantized electric field operator [23]. A contribution of the magnetic field can be written similarly. Here
$\displaystyle\mathcal{W}({\bf r},{\bf k},t)=\frac{1}{4\pi}\sum\limits_{\lambda_{\gamma},\tilde{\lambda}_{\gamma}}\int\frac{d^{3}{\tilde{k}}}{(2\pi)^{3}}\,\bm{E}^{*}_{\tilde{\lambda}_{\gamma}}({\bf k}-{\tilde{\bf k}}/2)\cdot\bm{E}_{\lambda_{\gamma}}({\bf k}+{\tilde{\bf k}}/2)\,e^{-it\left(\omega({\bf k}+{\tilde{\bf k}}/2)-\omega({\bf k}-{\tilde{\bf k}}/2)\right)+i{\bf r}\cdot{\tilde{\bf k}}}$ (3)
is a Wigner function of the photon, $\bm{E}_{\lambda_{\gamma}}({\bf k})=\frac{i\omega\sqrt{4\pi}}{\sqrt{2\omega n^{2}}}\,{\bm{e}}_{{\bf k}\lambda_{\gamma}}\sum\limits_{\lambda}\int\frac{d^{3}p}{(2\pi)^{3}}\,f_{e}^{(\text{in})}({\bf p},\lambda)S_{fi}^{(\text{pw})},\,{\bm{e}}_{{\bf k}\lambda_{\gamma}}\cdot{\bf k}=0,$ is a positive-frequency component of its electric field.
According to the standard interpretation [23], the energy density (2) defines probability to detect a photon in a region of space centered at the point ${\bf r}$ at the moment of time $t$. Clearly, the second marginal distribution $\int d^{3}x\,\mathcal{W}({\bf r},{\bf k},t)\propto|S_{fi}^{(\text{pw})}|^{2}$ yields probability to detect a photon with the frequency $\omega$ and a wave vector ${\bf k},\,|{\bf k}|=n(\omega)\,\omega$, the standard result of the quantum theory in momentum space (see [17]). Therefore, it is this Wigner function (3) in phase space that contains all the information on spatial distribution of the photon field at a given distance ${\bf r}$ – also in the near-field zone – and on its dynamics. Consequently, emission takes place in the pre-wave zone even if the condition of ChR is not met, but the waves do not constructively interfere to form a cone in the far field.
We calculate the Wigner function in the paraxial approximation, $\sigma\ll m$, in a medium with weak dispersion, $\frac{\omega}{n(\omega)}\frac{dn(\omega)}{d\omega}\ll 1$, and represent the amplitude [22, 17] as follows:
$S_{fi}^{(\text{pw})}=|S_{fi}^{(\text{pw})}|\,e^{i\zeta_{fi}}.$ (4)
Here $|S_{fi}^{(\text{pw})}|^{2}$ defines the emission rate in momentum space where the phase $\zeta_{fi}$ does not contribute to the probability, although it is non-vanishing even in the lowest order of the perturbation theory (Sec.5 in Supplementary). The result of the calculations is (Sec.3 in Supplementary)
[…]
where $G(t^{\prime})>0$ is not important for now, the momentum conservation ${\bf p}={\bf p}^{\prime}+{\bf k}$ is implied, and
$\displaystyle{\bm{R}}={\bf r}-{\bm{u}}_{p}t+(\partial_{{\bf p}}+\partial_{{\bf k}})\zeta_{fi}({\bf p},\lambda_{e},{\bf k},\lambda_{\gamma}).$ (6)
Here ${\bm{u}}_{p}={\bf p}/\varepsilon({\bf p}),\,\varepsilon({\bf p})=\sqrt{{\bf p}^{2}+m^{2}},\,{\bm{u}}_{k}={\bf k}/(n|{\bf k}|),|{\bm{u}}_{k}|=1/n$, $F(t^{\prime})\propto\arctan t^{\prime}/t_{d}$ contains a Gouy phase of the photon connected to its spreading with time $t^{\prime}$, and $t_{d}$ is a diffraction time (see below).
The principal behavior of the Wigner function (5), which is not everywhere positive, is governed by the ratio
$\displaystyle\frac{{\displaystyle R^{2}}}{{\displaystyle R_{\text{eff}}^{2}(t^{\prime})}}=\frac{{\displaystyle 1}}{{\displaystyle\sigma_{x}^{2}(t^{\prime})}}\Bigg{(}\underbrace{\frac{{\displaystyle[{\bm{R}}\times({\bm{u}}_{p}-{\bm{u}}_{k})]^{2}}}{{\displaystyle({\bm{u}}_{p}-{\bm{u}}_{k})^{2}}}}_{\text{finite at $t^{\prime}=0$}}+\text{terms vanishing at}\,t^{\prime}=0\Bigg{)},$ (7)
where $\sigma_{x}^{2}(t^{\prime})=\sigma^{-2}\left(1+(t^{\prime}/t_{d})^{2}\right)$ is an rms size of the electron packet. When the condition of ChR is met, $u_{p}>u_{k}$, the vector ${\bm{u}}_{k}-{\bm{u}}_{p}$ is directed backwards with respect to the electron velocity ${\bm{u}}_{p}$, and dependence of the Wigner function on ${\bm{R}}$ at small $t^{\prime}$ vanishes along ${\bm{u}}_{k}-{\bm{u}}_{p}$, defining the Mach cone with an angle
$\theta_{\text{Mach}}=\pi-\arcsin\left(\frac{{\displaystyle\sin\theta}}{{\displaystyle n|{\bm{u}}_{k}-{\bm{u}}_{p}|}}\right).$ (8)
If the electron is detected in a plane-wave state, scattered at the angles $\theta^{\prime},\phi^{\prime}$, the radius $R_{\text{eff}}(t^{\prime})$ depends on the difference $\phi_{R}-\phi$ between the azimuthal angle of ${\bm{R}}$ and that of ${\bf k}$, so it is anisotropic. The azimuthal symmetry of the Mach cone is restored when the electron is not detected and we integrate Eq.(5) over ${\bf p}^{\prime}$.
Figure 2: The effective correlation radius from Eq.(7) divided by the distance $u_{p}t^{\prime}$ traveled by the electron during the time $t^{\prime}$. Left: $\beta=0.7\,(\gamma=1.4),n=1.5,\,\theta_{\text{Ch.cl.}}=\arccos 1/u_{p}n\approx 17.8$ deg, $\theta_{\text{Mach}}\approx 107.8$ deg. Right: $\beta=0.9999\,(\gamma=70.7),n=1.33,\,\theta_{\text{Ch.cl.}}\approx 41.2$ deg, $\theta_{\text{Mach}}\approx 131.2$ deg. Nearby the Mach angle, space-time dependence of the Wigner function quickly vanishes within the correlation radius $R<R_{\text{eff}}(t^{\prime})$.
Figure 3: The spreading time (9) of the photon field measured in picoseconds (a. and d.) and its inverse (b.). The sharp maxima are nearby the angle $\theta_{\text{Ch.cl.}}=\arccos 1/u_{p}n$. The Cherenkov condition is not met for the black line in a.), which is why the photon field quickly spreads during hundreds of attoseconds. The right upper panel shows two points (10), between which the spreading time turns negative due to the quantum recoil ($\omega/\varepsilon\neq 0$, see Eq.(11)).
At large $t^{\prime}$, we find $R_{\text{eff}}^{2}(t^{\prime})\propto\sigma^{2}_{x}(0)\,t^{\prime 2}/t_{d}^{2}$ and when $R_{\text{eff}}(t^{\prime})\gg R$ dependence of the Wigner function on ${\bf r}$ and $t$ vanishes, whereas in the other limiting case, $R\gg R_{\text{eff}}(t^{\prime})$, it is exponentially suppressed. So, an effective region of space-time correlation is when $t^{\prime}\lesssim t_{d}$, and $R\sim R_{\text{eff}}(t^{\prime})$ is where it is most pronounced, which is why one can call $R_{\text{eff}}(t^{\prime})$ the correlation radius. At $t^{\prime}\gg t_{d}$, both the electron packet and the photon field spread and so there is no longer space-time correlation within the region $R<R_{\text{eff}}(t^{\prime})$, which is a hallmark of the wave zone. We show in Fig.2 that the correlation radius is orders of magnitude smaller than the distance $u_{p}t^{\prime}$ traveled by the electron for all the angles $\theta_{R}$ of ${\bm{R}}$, except for the Mach angle, $\theta_{R}\approx\theta_{\text{Mach}}$ (shown in Fig.3 c.)).
Spreading time and formation length
Let us discuss now the spreading time, which is found as
$\displaystyle t_{d}=\frac{2}{\sigma^{2}}\frac{({\bm{u}}_{p}-{\bm{u}}_{k})^{2}}{\left(1/\omega n^{2}-\varepsilon^{-1}\right)({\bm{u}}_{p}-{\bm{u}}_{k})^{2}+\left(\varepsilon^{-1}-\omega^{-1}\right)\left[{\bm{u}}_{p}\times{\bm{u}}_{k}\right]^{2}}.$ (9)
where $({\bm{u}}_{p}-{\bm{u}}_{k})^{2}=n^{-2}+u_{p}^{2}-2u_{p}\cos\theta/n,\ \left[{\bm{u}}_{p}\times{\bm{u}}_{k}\right]^{2}=u_{p}^{2}\sin^{2}\theta/n^{2}$ with the $z$ axis directed along the electron partial momentum ${\bf p}$. One can also define the formation length of radiation as $L_{f}=u_{p}|t_{d}|$, which turns to infinity at the Cherenkov angle in the classical Tamm problem [24, 10, 11, 25].
The diffraction time and the formation length have an extremum either at the angle $\cos\theta_{\text{Ch.cl.}}=1/u_{p}n<1$, when the Cherenkov condition $u_{p}>1/n$ is met, or at $\cos\theta=u_{p}n<1$ otherwise. Importantly, the Wigner function and the energy density stay finite in the latter case, but the photon field rapidly spreads (see the black line in Fig.3 a.)). Along with the extremum, the time $t_{d}$ has two points in which its denominator vanishes (see Fig.3 b.)),
$\displaystyle\cos\theta_{\infty}\approx\frac{{\displaystyle 1}}{{\displaystyle u_{p}n}}\left(1\mp\sqrt{\frac{{\displaystyle\omega}}{{\displaystyle\varepsilon}}}\sqrt{(n^{2}-1)(u_{p}^{2}n^{2}-1)}\right),$ (10)
where we have kept the first correction due to quantum recoil, which is usually very small [17], $\omega/\varepsilon\ll 1$. Clearly, the diffraction time can only turn to infinity under the condition of ChR, $u_{p}n>1$. The angular width between the two points is
$\Delta\theta_{\infty}\approx 2\sqrt{\frac{{\displaystyle\omega}}{{\displaystyle\varepsilon}}}\sqrt{n^{2}-1},$ (11)
and it vanishes for classical emission with no recoil, $\omega/\varepsilon\to 0$. For materials like Al, Si, Be, and Ti, Cherenkov radiation can be observed in the soft X-ray range at the frequencies [26, 27, 25] $\omega\approx 72.5$, $100$, $110$, $453.8$ eV, respectively, which for $\varepsilon\sim(5-20)m$ yields $\Delta\theta_{\infty}<1-2\,\text{deg}$ (cf. Fig.3 c.), d.)).
Between the above points, the time $t_{d}$ becomes negative – see Fig.3 b.) – and the Gouy phase $\arctan t^{\prime}/t_{d}$ changes it sign, as if the electron packet shrinks during the emission. In the classical regime with no recoil, both the points merge and so $t_{d}$ and $L_{f}$ turn to infinity at the Cherenkov angle. Indeed, in a vicinity of this angle the spreading time with the recoil kept is
$\displaystyle t_{d}\Big{|}_{\cos\theta=1/nu_{p}}=\frac{2\varepsilon}{\sigma^{2}}\frac{n^{2}}{1-n^{2}}<0,$ (12)
where $\varepsilon=\gamma m$. We compare this with the spreading time of an electron packet, which is at rest on average in vacuum [28] $t_{d}^{(\text{e, rest})}=m/\sigma^{2}$. In the laboratory frame this time is $\gamma$ times larger, which coincides with $|t_{d}|$ up to the factor $22n^{2}/(n^{2}-1)>$ 2. So, in a vicinity of the Cherenkov direction spreading of the photon seems to reverse back and it is intimately connected with spreading of the electron packet.
[…]
Shift of the photon arrival time and flash duration
Dependence of the Wigner function (5) on the detection time $t$ comes exclusively from the following envelope:
$\displaystyle\exp\left\{-\frac{{\displaystyle R^{2}}}{{\displaystyle R_{\text{eff}}^{2}(t^{\prime})}}\right\}\propto\exp\left\{-\frac{(t-t_{0})^{2}}{2\sigma_{t}^{2}(t^{\prime})}\right\},$ (13)
$\displaystyle\sigma_{t}^{2}(t^{\prime})=\frac{\sigma^{2}_{x}(t^{\prime})}{2}\frac{({\bm{u}}_{p}-{\bm{u}}_{k})^{2}}{\left[{\bm{u}}_{p}\times{\bm{u}}_{k}\right]^{2}}.$ (14)
Here natural duration of the Cherenkov flash is defined by $\sigma_{t}(t^{\prime})$ and a time instant at which the probability to catch the photon around the point ${\bf r}$ is maximized is $t_{0}={\bm{l}}_{0}\cdot\left({\bf r}+(\partial_{{\bf p}}+\partial_{{\bf k}})\zeta_{fi}\right),\,{\bm{l}}_{0}=\left[({\bm{u}}_{p}-{\bm{u}}_{k})\times\left[{\bm{u}}_{k}\times{\bm{u}}_{p}\right]\right]/\left[{\bm{u}}_{p}\times{\bm{u}}_{k}\right]^{2}$.
One can neglect the term with the phase $\zeta_{fi}$ in ${\bm{R}}$ (6) in the wave zone where $r\to\infty$, and then the Wigner function (5) and the emitted energy are concentrated in a vicinity of the electron classical trajectory [34], ${\bf r}\sim{\bm{u}}_{p}t$. The detector registers a photon in the far field emitted at $t=0,r_{0}=0$ by a classical point-like electron at the time instant
$t_{\text{cl.}}^{(\text{far-f.})}=r/u_{k}=r\,n,$ (15)
which will be called the classical arrival time. Let us compare this prediction with the above $t_{0}$, derived quantum mechanically. Directing the $z$ axis along the electron momentum ${\bf p}$, we find ${\bf k}=n\omega{\bm{l}},\ {\bm{l}}=\{\sin\theta\cos\phi,\sin\theta\sin\phi,\cos\theta\},\ r\equiv{\bf r}\cdot{\bm{l}},\ t_{\text{cl.}}^{(\text{far-f.})}={\bf r}\cdot{\bm{l}}\,n,$ and ${\bm{l}}_{0}=\sin^{-2}\theta\,\left(\left(u_{p}^{-1}-n\cos{\theta}\right){\bf p}/|{\bf p}|+\left(n-u_{p}^{-1}\cos\theta\right){\bm{l}}\right)$. In a vicinity of the Cherenkov angle, we have ${\bm{l}}_{0}\to{\bm{l}}\,n$ and so for $\zeta_{fi}=0$ we get $t_{0}\to t_{\text{cl.}}^{(\text{far-f.})}={\bf r}\cdot{\bm{l}}\,n$, in accord with Eq.(15).
However the phase $\zeta_{fi}$ cannot be ignored in the formation zone and this affects the photon arrival time $t_{0}$. In this regime, we call $t_{\text{cl.}}={\bf r}\cdot{\bm{l}}_{0}$ for arbitrary emission angles and so the quantum shift is
$\Delta t=t_{0}-t_{\text{cl.}}={\bm{l}}_{0}\cdot(\partial_{{\bf p}}+\partial_{{\bf k}})\zeta_{fi}.$ (16)
The physical origin of this delay is the electric dipole moment density $\propto e(\partial_{{\bf p}}+\partial_{{\bf k}})\zeta_{fi}$ induced in medium by the field of the electron. We deal with an analogue of the AC Stark effect [15] with the atoms being off-resonantly polarized by a broadband spectrum $\Delta\omega$ of pseudo-photons. Similarly to the observed time delays when a laser propagates in a medium [35, 36, 37], here we encounter delays induced by the virtual photons, reemitted as real ones. Classically, one can look at this as if the photon was emitted not from a point-like electron, but from a point shifted laterally to $\Delta\rho\sim\beta\gamma\lambda/2\pi$ from its trajectory, which is a mean free path of the virtual photon [20, 11]. Numerically $\Delta\rho/c\sim\beta\gamma\lambda/2\pi c=\beta\gamma/\omega\sim 1\,\text{fs}-100\,\text{fs}$ for photons from IR to UV ranges and $\gamma=\varepsilon/m\sim 10$.
Figure 4: The quantum shift $\Delta t=t_{0}-t_{\text{cl.}}={\bm{l}}_{0}\cdot(\partial_{{\bf p}}+\partial_{{\bf k}})\zeta_{fi}$ from Eq.(16) of the photon arrival time compared to the classical value and the Cherenkov flash duration (the green lines with $\sigma=p_{\perp}$). The electron energy is typical for a transmission electron microscope: $\beta=0.7$, and $p^{\prime}_{\perp}=0.99\times p_{\perp},p^{\prime}_{z}=0.9\times\beta,n=1.5$. The left panel (a.): $p_{\perp}=10^{-5}\,m,\,1/p_{\perp}\gtrsim 10\,\text{nm}$, the right panel (b.): $p_{\perp}=10^{-6}\,m,\,1/p_{\perp}\gtrsim 100\,\text{nm}$. The behaviour at small angles is shown in the insets. The shifts vanish in those regions that are not allowed by the momentum conservation law, they stay roughly the same for other $p^{\prime}_{\perp},p^{\prime}_{z}$ and for ultrarelativistic electrons, $\gamma\gg 1$, though the Cherenkov angle grows.
Figure 5: Dependence of the shift $\Delta t=t_{0}-t_{\text{cl.}}={\bm{l}}_{0}\cdot(\partial_{{\bf p}}+\partial_{{\bf k}})\zeta_{fi}$ from Eq.(16) of the photon arrival time on the refractive index $n$ at the emission angle $\theta=10$ deg. Left: $\beta=0.7,p^{\prime}_{z}=0.9\times\beta$ (a TEM regime; the minimal refractive index for which the Cherenkov condition is met is $n\approx 1.429$, whereas $n=1.45-1.55$ for fused silica in the given spectral range), right: $\beta=0.9999(\gamma=70.7),p^{\prime}_{z}=0.99\times\beta$ (an accelerator regime). The classical Cherenkov angle $\theta_{\text{Ch.cl.}}=\arccos 1/\beta n$ is shown in the insets.
When measuring the quantum shift in the photon arrival time from its classical value, the flash duration $\sigma_{t}(t^{\prime})$ is crucial because the deviations can hardly be discerned when $\sigma_{t}(t^{\prime})\gg|\Delta t|$ occurring for $t^{\prime}\gg t_{d}$ far from the Cherenkov angle. This duration was estimated by Frank as early as in 1956 from the classical considerations [8, 9] to be $\sim 1/\Delta\omega<1\,\text{ps}$ where $\Delta\omega$ is a frequency interval for which the emission takes place. The quantum estimates from the uncertainty principle for the off-resonant AC Stark effect yield the same result [15], $1/\Delta\omega\sim 0.1-100\,\text{fs}$ for $\Delta\omega\sim 10^{-2}-10$ eV. Our quantum model predicts the following flash duration nearby the Cherenkov angle:
$\sigma_{t}(t^{\prime})\to n\,\sigma_{x}(0)/\sqrt{2},$ (17)
because in this case $t_{d}\to\infty$. Clearly, only the electron packet’s size at the target entrance ($t^{\prime}=0$) contributes to the flash duration nearby $\theta_{\text{Ch.cl.}}$. For realistic electrons with $\sigma_{x}(0)\sim 1-100\,\text{nm}$, we find
$\sigma_{t}(0)\sim 10\,\text{as}-1\,\text{fs}.$ (18)
In Figures 4 and 5 we show that the typical shifts indeed belong to the attosecond range and that the flash duration is larger than the shift, except for small angles. The electron transverse momenta are chosen to be $p_{\perp}=\sigma\sim(10^{-7}-10^{-4})\,m$ because they correspond to the spatial widths $1/p_{\perp}=\sigma_{x}(0)\gtrsim 1\,\text{nm}-1\,\mu\text{m}$, respectively, and we also neglect the spreading. Note that within the pre-wave zone the emission does not take place only at the Cherenkov angle. The sign of the shift swaps between the two kinematic scenarios (Sec.4,5 in Supplementary), which is why the absolute value $|\Delta t|$ is shown. Fixing the detector at certain angles $\theta,\phi$ and the distance $r$, one would see that the photons equally probably arrive either later (time delay) than $t_{\text{cl.}}$ or sooner than that (negative delay), if the electron is not detected. Note that integration of the Wigner function over ${\bf p}^{\prime}$ puts the electron momentum to the definite value ${\bf p}^{\prime}={\bf p}-{\bf k}$ with subsequent integration over ${\bf p}$ with the Gaussian distribution. If one wishes to catch only the shifts with one particular sign, one should detect the photon and the electron in coincidence, which is technically more challenging.
In Figure 5 we highlight that the flash duration $\sigma_{t}$, shown with the green line, can drop below the temporal shift for angles $\theta\lesssim 10-15$ deg and media with small refractive indices $n\sim 1.01-1.5$, however, only with no spreading taken into account as $\sigma_{t}(t^{\prime})$ grows with the time $t^{\prime}$, whereas the shift does not. For the energies of transmission electron microscopes (TEM), $\beta\sim 0.7$, a target made of fused silica with $n=1.458$ in the optical range can do the job, whereas for ultrarelativistic electrons, $\gamma\gg 1$, aerogels with $n\sim 1.01-1.3$ can be employed, which are already used as Cherenkov generators [38, 39, 40]. The use of Brewster-Cherenkov detectors [41] can also come in handy here.
If we now go beyond the model of the electron packet with $\sigma_{\perp}=\sigma_{||}$ and recall that there is Lorentz contraction, $\sigma_{||}=\gamma^{-1}\sigma_{\perp}$, the flash durations can become $\gamma$ times shorter than (18). Although photon spreading can be safely neglected nearby the Cherenkov angle, spreading of the electron before entering the target can significantly increase these numbers. According to the quantum dynamics of the coherence length $\sigma_{x}(t^{\prime})$ – also called the generalized van Cittert-Zernike theorem [28] – when an electron is released from a photo-gun or a field emitter with rms sizes $\sigma_{\perp}\sim\sigma_{||}$ [29] of a few nanometers and an energy up to a few tens of eV, it spreads to some tens of nanometers at the distance of 1 $\mu\text{m}$ and it reaches micrometer sizes already at 1 mm. Clearly, electron acceleration – say, in an RF cavity – mitigates the spreading rate, but a realistic estimate of the packet length $\sigma_{||}$ at the target boundary remains unknown.
In practice, achieving attosecond flash durations requires electrons generating radiation with no spreading before the target, just after the emission from a cathode. A vacuum gap of even a few millimeters after the source would yield micrometer-sized electron packets entering the target, which results in picosecond durations (subpicosecond with acceleration to a few MeV) measurable by streak cameras. State-of-the-art Cherenkov counters at accelerators have picosecond time resolution [40, 42], the subfemtosecond resolution can be achieved at X-ray free-electron lasers [43, 44] or with the frequency-resolved optical gating [45], whereas it is nanoseconds for Cherenkov telescopes employed in gamma-ray astronomy [46] where spreading effects are notable. Attosecond photon pulses – including twisted photons with orbital angular momentum – are usually obtained through high-harmonic generation in the extreme-ultraviolet and even soft X-ray ranges [47, 48], enabling vortex electrons generation via photoionization [48]. The ChR can be a source of twisted photons [17, 49], also in the soft X-ray range, and shaping the spatial and temporal profiles of the electron wave function offers refined tuning of phase-space profile of the photons.
Close to the Cherenkov angle, the flash duration is directly tied to the electron coherence length upon entering the medium, opening avenues for contolling this duration by selecting packets of desired length and accelerating them in the low-current regime with no space-charge effects, typical for TEMs. Alternatively, measuring photon pulse durations can provide a novel method for determining electron packet length, a complementary approach to interferometry that is reminiscent of bunch length measurements via classical coherence [13]. Naturally, ChR serves as an exemplary case within a broader class of media-induced emission phenomena where attosecond time scales are accessible via phase-space analysis.
Acknowledgment. We are grateful to D. Glazov, I. Ivanov, A. Shchepkin, V. Serbo, I. Pavlov, and especially to A. Tishchenko for useful discussions and suggestions, and also to Maria Zhuravleva for the help with the 3D picture. The studies on photon emission in phase space are supported by the Foundation for the Advancement of Theoretical Physics and Mathematics “BASIS”. Those on the temporal features of radiation at the finite distance by the Russian Science Foundation (Project No. 23-62-10026 https://rscf.ru/en/project/23-62-10026/). The analysis of the quantum shift is supported by the Government of the Russian Federation through the ITMO Fellowship and Professorship Program.
Author contributions. All authors contributed extensively to this work. Dm.K. conceived the idea, developed the theoretical framework, wrote the manuscript with input from all the authors and acquired the funding. D.G., G.S., A.C., and D.K. checked the calculations and interpreted the results in detail. D.G., A.C., and G.S. assisted with the figures, while A.C., Dm.K., and G.S. refined the presentation of results and theoretical details.